OpenStudy (anonymous):

The football team has 30 more members than the basketball teams. If each group had 10 more members, the ratio of their membership would be 3:2. How many members are in each group? Write your answer as football x basketball y

OpenStudy (anonymous):

let basketball team has x members than football team has x +30members if 10 more members, the ratio of their membership would be 3:2 that is x+30 +10 / (x +10) = 3/2 x+40/(x+10) = 3/2 2(X+40) = 3(x+10) 2x + 80 = 3x + 30 x = 50 basketball team has 50 members and football team has 50 + 30 = 80 members